OpenStudy (anonymous):

What is the equation of the line meeting the given conditions. Please put the equation in standard form. Containing E(4, 3) and F(6, 1)

OpenStudy (anonymous):

Use the formula for rise orver run and find the slope of the line, then use y=mx+b to find the y-intercept and you will have the formula

OpenStudy (anonymous):

so the slope is -1?

OpenStudy (anonymous):

Yes, now use y+mx+b to find the y intercept

OpenStudy (anonymous):

ok so what is b?

OpenStudy (anonymous):

3=(-1)4+b so, 3=(-4)+b 3-(-4)=b 7=b,

OpenStudy (anonymous):

ok but it needs to be in y-y1=m(x-x1) form

OpenStudy (anonymous):

Oh, um sorry :$ I forgot how to do that.. but i think you need the x-intercept... and its supposed to be like 7-7=(-1)({x-intercept}-{x-intercept})